A quality control inspector has drawn a sample of 18 light bulbs from a recent production lot. If the number of defective bulbs is 2 or less, the lot passes inspection. Suppose 30% of the bulbs in the lot are defective. What is the probability that the lot will pass inspection? Round your answer to four decimal places.
Answer #1

Answer:

Given,

sample n = 18

p = 0.3

q = 1 - 0.3

= 0.7

P(X <= 2) = P(0) + P(1) + P(2)

= 18C0*0.3^0*0.7^18 + 18C1*0.3^1*0.7^17 + 18C2*0.3^2*0.7^16

= 0.0016 + 0.0126 + 0.0458

= 0.0600